A building contractor hires 8 bricklayers and 2 roofers, his daily payroll is #960, while 10 bricklayers and 5 roofers require a daily payroll of #1500. What is the daily payroll of a bricklayer and a roofer?
Answer by Theo(13342)

x = number of bricklayers
y = number of roofers

8x + 2y = 960
10x + 5y = 1500

multiply both sides of the first equation by 5 and multiply both sides of the second equation by 2 to get:

40x + 10y = 4800
20x + 10y = 3000

subtract the second equation from the first to get:
20x = 1800

solve for x to get:
x = 90

replace x in the first original equation by 90 to get:
8x + 2y = 960 becomes 8 * 90 + 2y = 960
simplify to get:
720 + 2y = 960
subtract 720 from both sides to get:
2y = 240
solve for y to get:
y = 240/2 = 120

you have:
x = 90
y = 120

replace x and y in both original equations to get:

8x + 2y = 960 becomes 8 * 90 + 2 * 120 = 960 which becomes 720 + 240 = 960 which becomes 960 = 960.

10x + 5y = 1500 becomes 10 * 90 + 5 * 120 = 1500 which becomes 900 + 600 = 1500 which becomes 1500 = 1500

this confirms the solution is correct.

the solution is:

the daily payroll of a bricklayer is 90.
the daily payroll of a roofer is 120.